OpenStudy (henrietepurina):

Identify the vertex for the graph of y = 2x^2 + 8x - 3.

OpenStudy (henrietepurina):

@ganeshie8

OpenStudy (triciaal):

general form ax^2 + bx + c = 0 vertex form x-coordinate = -b/2a y coordinate = f(x)

OpenStudy (triciaal):